eferences Mailings Review ViewTable Design Layout Question XYZ restaurant sells donuts at $2 per unit. It uses capital (which it rents at $5 per hour under a contract for 5 and labor (which is paid a wage of $150 per hour of labor services). Complete the following table and use that information to answer the following questions. VC MPL APL VML FC 100 25 50 75 50 150 25-50 125 125100 625 200 2 300 125 75 250 450 225 600 150 25 93.75 1 75 375 750 25 450 450 150 175 1100-1900 400 25574 150 25 1050 57.14 150 1200 100 25 1350 100 9 450 5050 1500 10 500 25SO MacBook Ai
NG 640 VVK 6 Hw (1) ut References Mailings Review View Table DesignLayout e. What is the variable cost of producing 375 units of output? WHY f. How many units of the variable input should be used to maximize profits? SHOW WORK g. Define profit? What are the maximum profits this firm can earn? SHOW WORK Question 2 There is no difference between the law of diminishing marginal returns and the law of diminishing marginal rate of technical substitution. True or False. Explain and offer examples to further illustrate your explanation. Home Assignment/Drop Box Instructions
Answer #1

e. Variable cost at 375 units is 600.

( because, to produce 375 units in 1 hours, 4 laborers are hired at $150/hour. So, 150*4 = 600 )

f. The number of labor units required to maximize profits can be identified by observing the marginal productivity of labor (MPL) and the quantity produced

As the units of labor increased at first it increased , then started to reduce and became negative. The labor unit at which there is no increase in productivity that is MPL is 0 , is the number of variable units required to maximize profits.

MPL 150 100 50 2 10 12 50 -100 LABOR dW

In this case at 5 units of labor , the MPL is 0 , if the firm hires more laborers then the firm will have suffer losses. Before 5 , the firm has the ability to increase labor units to produce more units and earn higher revenue.

So, at 5 units maximum profits will be attained.

g. Profit is the difference between the revenue earned from selling the final good and the costs incurred to produce the good. It is the net benefit for the firm.

Profit = Revenue- Cost

As mentioned in the previous answer, at 5 units profits will be maximized because before that the firm has the ability to earn more and more labor will reduce the profits and maybe the firm will incur a loss.

at, 3 units

Revenue = 225*2 - 450

Costs = FC + VC = 25 + 450 = 475

Loss = 25

So, at 4 units of labor :

Revenue = 375 * 2 = 750

Costs = FC + VC = 25 + 600 = 625

Profit = 750 - 625 = 125

at 5 units of labor :

Revenue = 450*2 = 900

Cost = FC + VC = 25 + 750 = 775

Profit = 900 - 775 = 125

at 6 units of labor :

Revenue = 450 * 2 = 900

Cost = FC + VC = 25+900 = 925

Loss = 25

Maximum profit it can earn is $125 at 5 units of labor

2)

False. Both the laws are pertaining to different concepts.

The law of diminishing marginal returns refers to how employment of additional variable factor leads to reduction in output.

where as, The law Diminishing Marginal Rate of Substitution states that a consumer who has 2 goods X and Y, as the consumption of X increases the consumer is willing to giving up less and less of good Y.

For (e.g )

law of diminishing marginal returns

L MP
1 2
2 4
3 6
4 8
5 10
6 8
7 5

MP 12 10 8 6 2 0 4 6 8 LABOR LC dW

You, can observe that till the 5 unit of labor, the marginal productivity is increasing but beyond that it starts to fall.

So, after a certain point , additional unit of labor employed will lead to diminished output/returns.

Law of Diminishing Marginal Rate of Substitution

X Y MRSyx
1 12 4
2 8 3
3 5 2
4 3 1
5 2

It can be observed that at first the consumer is willing to give up 4 units of y for one unit of X which consequently keeps on reducing. So, the graph is convex to the origin

Y 14 12 10 8 6 2 2 3 6 X Ln t